OpenStudy (anonymous):

A. Solve the equation for y in terms of x, and replace y with the function notation f(x). B. Then find f(0) X+4y=11 I think I got it, I'm just not sure which one goes to which, a or b. I got 11-x/4 and x=11

OpenStudy (anonymous):

first part is right

OpenStudy (anonymous):

\[f(x)=\frac{11-x}{4}\]

OpenStudy (anonymous):

second part is not right you solved \[\frac{11-x}{4}=0\] and got \(x=11\) but it asked for \[f(0)\]

OpenStudy (anonymous):

so if \[f(x)=\frac{11-x}{4}\] then \[f(0)=\frac{11-0}{4}\]

OpenStudy (anonymous):

Oh okay, that make sense. Thank you so much!

OpenStudy (anonymous):

yw